OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = −x + 7

OpenStudy (paxpolaris):

\[-{1 \over 2 x + 2} = −x + 7\]\[\implies {1 \over 2 x + 2} = x - 7\]\[\implies 1=(x-7)(2x+2)=\large 2x^2-12x-14\] \[\Large \implies 2x^2-12x-15=0\]

OpenStudy (paxpolaris):

\[x={6\pm \sqrt{36-30}\over2} \huge =3\pm {\sqrt 6 \over 2}\]